OpenStudy (anonymous):

Y is the midpoint of segment XZ. The coordinates of X are (2, 4) and of Y are (-1, 1). Find the coordinates of Z.

OpenStudy (anonymous):

To find midpoint it is the average distance between two different endpoints. So since midpoint is Y. X(2,4) Z(Z1,Z2) ... (2+Z1)/2 = -1 and (4+Z2)/2 = 1. Solve for Z1 and Z2. An ambiguous question: If Y is midpoint of XZ. X(X1,X2) Z(Z1,Z2) Y(Y1,Y2) (X1+Z1)/2 =Y1 (X2+Z2)/2= Y2